# Intrinsic Dimension Correlation: uncovering nonlinear connections in multimodal representations

### v1

@@ 3.1 Synthetic experiments
As a first step, we evaluate our correlation method in a simple and controlled setting. We produce three 2D datasets, displayed in the Appendix in Fig. 8. Such datasets are made of two variables that are linearly correlated, uncorrelated, and nonlinearly correlated. The first setting is simply obtained by arranging $x$ and $y$ on a straight line, in the second case both random variables are sampled independently from a normal Gaussian distribution, while the last dataset contains data arranged on a spiral curve.
@@
We report our correlation results in Table 1. In the simpler cases (linear and random data) our method agrees with linear correlation and distance correlation, correctly identifying a very strong correlation in linear data and the lack thereof in random data. The spiral dataset constitutes a more tricky testbed: while there is a clear correlation between the two variables, it is a highly nonlinear one, and the linear correlation coefficient is around 00. Even distance correlation, despite being a nonlinear method, fails to find any strong signal of correlation, returning a value very close to 00. Instead, our method correctly identifies the strong dependency between the two variables, with a correlation coefficient of $0.94$, determined with high confidence, as witnessed by the $p$-value of $0.01$. We note here that our method relies on an intrinsic dimension estimator (in this paper, we employ TwoNN [27]), and it inherits substantial properties from it. On the negative side, $I_{d}$ estimators are not oracles, and they can return values that slightly differ from what one would expect (e.g. $I_{d}$ lower than $1$ in our linear dataset), or even totally fail (when the $I_{d}$ becomes large enough, the estimator is also affected by the curse of dimensionality). Conversely, the choice of employing TwoNN makes our method extremely efficient (the correlation coefficient can be obtained with just 3 calls to the estimator), allowing it to scale easily to large and high-dimensional datasets. Moreover, it is worth noting that, since TwoNN only relies on local distances, up to the second nearest neighbor, our correlation method is automatically invariant to any transformation that preserves the local neighborhood structure of data points.
Table 1: Correlation in 2D datasets. ($\mathbf{R^{2}}$): linear correlation coefficient; ($\mathbf{dCor}$): distance correlation coefficient; ($\mathbf{I_{d}}$): intrinsic dimension of the concatenated (2D) dataset; ($\mathbf{I_{d}Cor}$): correlation coefficient ($\varrho$) returned by our method; ($\mathbf{p}$-value): significance of the correlation detected by our method ($100$ shuffles).
Data $\mathbf{R^{2}}$ $\mathbf{dCor}$ $\mathbf{I_{d}}$ $\mathbf{I_{d}Cor}$ $\mathbf{p}$-value
Linear $1.00$ $1.00$ $0.98$ $1.00$ $0.01$
Random $0.00$ $0.00$ $1.99$ $0.00$ $0.37$
Spiral $0.01$ $0.02$ $1.07$ $0.94$ $0.01$
4 Results
In this section, we assess our proposed $I_{d}$or algorithm using representations generated by different neural networks on various datasets. We begin with a straightforward example that underscores the difficulty faced by standard methods in identifying nonlinear relationships, then progress to more extensive applications involving visual representations and multimodal text-image representations. For some experiments, we do not report the correlation $p$-value returned by our method. In such cases, the $p$-value is always the lowest possible according to the formula in Algorithm 1 ($\frac{1}{S+1}$, where $S$ is the number of permutations, $100$ in most of our experiments).
4.1 A motivating example
@@
Figure 2: Correlation results with different methods between MNIST data and their final representations computed by a randomly initialized MLP, with variable degree of activation nonlinearity, increasing on the $x$ axis.
@@
To illustrate this phenomenon, we showcase a simple example: we consider a randomly initialized multilayer perceptron (MLP), made of $15$ fully connected layers of $784$ neurons, followed by a LeakyReLU activation. LeakyReLU is a parametric activation function, whose behavior depends on a parameter called slope: if the slope is $1$, it behaves like the identity function, rendering our MLP a linear function of the input, while lower slope values make the network nonlinear, with 00 corresponding to the standard ReLU. We feed our MLP with the MNIST [39] dataset at increasing degrees of nonlinearity (which corresponds to decreasing the slope) and compute the correlation between the representation at the final layer and the input data, both with our method and with established baselines (SVCCA, Distance Correlation, linear kernel CKA, and RBF kernel CKA).
@@ 4.2 ImageNet representations
Moving to a more realistic setting, we test our method on measuring similarity between ImageNet [40] embeddings coming from different neural encoders. We consider a variety of pre-trained architectures, including CNNs (ResNet-18 [41] and EfficientNet-B0 [42]), four variants of Vision Transformers [43] (including a ViT-CNN hybrid) and two ViT-based models trained on a self-supervised vision-language loss (CLIP-ViT-B [6] and SigLIP-ViT-B [44]). For all the models, we consider the output of the last hidden representation produced by the encoder, before the classification head: in supervised ViTs this choice corresponds to the last class token, while in CNNs to the pooler output. For CLIP and SigLIP, we consider instead the image representation in the shared vision-language space.
Figure 3: Correlation results on ImageNet representations, obtained using: left $I_{d}$or (ours); right Distance Correlation. Both methods are able to detect non-negligible correlation. More baseline results are reported in the Appendix.
@@
We report the correlation results produced by our algorithm in Fig. 3, along with the correlation scores returned by Distance Correlation (dCor), as it is the baseline method that most closely matches ours in terms of mean off-diagonal correlation ($I_{d}$or mean: $0.87$, dCor mean: $0.51$). Detailed results for SVCCA (mean: $0.44$), linear kernel CKA ($0.43$) and RBF kernel CKA ($0.44$) are reported in the Appendix in Fig. 6. In this setting, all models are computing embeddings for the same data points, hence we would expect significant correlation to be present for any given pair of models. Indeed, all methods are clearly capturing such similarity, including SVCCA, which suggests that a relevant component of this correlation is actually linear.
@@ 4.2.1 Coarse alignment
The previous section demonstrated that our method effectively identifies strong correlations within perfectly aligned datasets of representations. We now aim to explore how the method’s performance might vary when applied to coarsely aligned data. To test this, we utilize the inherent class information of ImageNet data. Specifically, we randomly shuffle the embeddings generated by a model while keeping the labels unchanged, and then compare this modified dataset with the original dataset prior to shuffling. In other words, given a point index $i$ of class $C_{i}$, we pair it with another randomly chosen point $j$ of class $C_{j}$ with the condition that $C_{i}=C_{j}$. This allows us to assess the robustness of our correlation estimation in less ideal conditions: we expect the correlation signal to decrease, as alignment is a crucial property for our method to detect similarities. However, as we report in Table 3, we are still able to identify correlations with high confidence, as demonstrated by the low $p$-values. Interestingly, the values of the correlation suggest that the number of features needed to perform the classification task is between 50 and 70 %. Moreover, we observe that the models that exhibit lower correlation are CLIP and SigLIP, which have been trained with a self-supervised loss, hence without the explicit notion of class. For reference, we also report the $p$-values obtained when one of the two datasets is freely shuffled (irrespective of class labels), which correctly reports a lack of correlation. It is also worth noting that, while the average value of the correlation for $I_{d}$or is $0.63$, the values for the baselines are in the range $(0.32-0.48)$.
Table 2: Correlation between ImageNet representations when exact alignment is broken. Results are shown, in terms of $p$-value and correlation coefficient, for the coarse alignment case, in which data are shuffled while preserving the labeling. The last row reports the $p$-value for the fully shuffled case, in which no constraint is enforced on labels, and correlation is not present.
Metric EfficientNet SigLIP ViT-B-16 ViT-B-32 ViT-hyb. ViT-L ResNet CLIP
$p$-val. (coarse) $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$
$I_{d}$or (coarse) $0.61$ $0.52$ $0.68$ $0.69$ $0.63$ $0.72$ $0.66$ $0.52$
$p$-val. (rand) $0.11$ $0.64$ $0.89$ $0.80$ $0.12$ $0.19$ $0.74$ $0.45$
@@ 4.3 Multimodal representations
We now shift our focus to a multimodal context, where we examine the similarities between hidden spaces learned by text and image encoders. We use two datasets, N24News [45] and MS-COCO 2014 [46], both of which consist of image-caption pairs. This analysis will help us understand how textual and visual representations correlate when evaluated on related multimodal content. Images are encoded using a representative subset of the vision models introduced in Sec. 4.2: two CNNs (EfficientNet-B0 and ResNet-18), two ViTs (ViT-B-16 and ViT-hybrid), and the visual branch of CLIP. For text we employ five architectures, all taken pretrained: BERT [47], both cased and uncased, ALBERT [48], Electra [49] and finally the text encoder of CLIP. For all text models, we consider the last representation of the class token.
We report the correlation results obtained on N24News in Fig. 4, comparing our method against Distance Correlation (dCor), which is once again the closest-performing baseline method. Our $I_{d}$or yields a mean off-diagonal correlation of $0.66$, which is noticeably higher than those of baseline methods, in the range $(0.25-0.29)$. In fact, the correlation heatmaps for all baseline methods reveal a clear block structure, as such methods are able to capture correlation only among same-modality encoders, but fail on cross-modal correlation. Full results for baseline methods are available in the Appendix (Fig. 7). Instead, our method returns significant correlation values even across different modalities, in accordance with previous findings [17] that proved N24News representations to be transferable across models and modalities. Experiments on MS-COCO confirm the behavior observed for N24News, as discussed in the Appendix (section A).
Figure 4: Correlation results on N24News representations, obtained using: left $I_{d}$or (ours); right Distance Correlation (dCor). Like all other baselines we evaluate, dCor is only able to spot correlations between encoders of the same modality, while $I_{d}$or reveals significant correlation for all model pairs.
@@ 4.4 Computational resources
We performed all the computations on a NVIDIA A100 GPU, equipped with 40GB of RAM. The main computational hurdle of our method is the estimation of $I_{d}$ through TwoNN: our implementation follows closely that of [28], which we translated to PyTorch to enable GPU acceleration. With this setup, $I_{d}$or runs in the order of $1$s on two $1024$-dimensional datasets of $30000$ points. Just like our correlation method, all the representation similarity baselines we employ greatly benefit from GPU acceleration: we used them in their PyTorch implementations provided by [50] (SVCCA), [51] (CKA) and [8] (Distance Correlation), with minor adaptations. Pretrained models were obtained from the Transformers library by HuggingFace [52], details on the checkpoints we employed are provided in the Appendix (section D).
@@ 5 Discussion
Our work introduces a novel and robust method for detecting complex nonlinear correlations in high-dimensional spaces. Due to its flexibility, this method can be employed in a wide range of applications, from natural language processing to computer vision and beyond (including other fields of science, like physics), offering a new type of analysis to address how machine learning models represent data. Remarkably, our results show the effectiveness of the method to detect a correlation signal in multimodal data where we know a correlation should exist but where standard methods struggle to identify any.

### v2

@@ 4.1 Synthetic experiments
As a first step, we produce three toy datasets, displayed in the Appendix in Fig. 5. Such datasets are made of 5000 observations of two variables that are either linearly correlated, uncorrelated, or nonlinearly correlated. The first setting is simply obtained by arranging $x$ and $y$ on a straight line, in the second case both random variables are sampled independently from a normal Gaussian distribution, while the last dataset contains data arranged on a spiral curve.
@@
We report our correlation results in Table 1, aggregating results over 10 independent random samplings of the datasets. In the simpler cases (linear and random data) our method agrees with linear correlation and distance correlation, correctly identifying a very strong correlation in linear data and the lack thereof in random data. The spiral dataset constitutes a more tricky testbed: while there is a clear correlation between the two variables, it is a highly nonlinear one, and the linear correlation coefficient is around 00. Even Distance Correlation, despite being a nonlinear method, fails to find any strong signal of correlation, returning a value very close to 00. Instead, our method correctly identifies the strong dependency between the two variables, with a mean correlation coefficient of $0.98$, determined with high confidence, as witnessed by the $p$-value consistently equal to $0.01$. We note here that our method relies on an intrinsic dimension estimator (in this paper, we mainly employ TwoNN (Facco et al., 2017), but we also provide results on synthetic data using MLE (Levina & Bickel, 2004) in section A.3.1 in the Appendix), and it inherits substantial properties from it. On the negative side, $I_{d}$ estimators are not oracles, and they can return values that slightly differ from what one would expect (e.g., $I_{d}$ lower than $1$ in our linear dataset or higher than $2$ in the random case), or even totally fail (when the $I_{d}$ becomes large enough, the estimator is also affected by the curse of dimensionality). Conversely, the choice of employing TwoNN makes our method extremely efficient (the correlation coefficient can be obtained with just 3 calls to the estimator), allowing it to scale easily to large and high-dimensional datasets.
[…]
Data $\mathbf{R^{2}}$ $\mathbf{dCor}$ $\mathbf{I_{d}\oplus}$ $\mathbf{I_{d}Cor}$ $\mathbf{p}$-value
Linear $1.00\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.02$ $1.00\pm 0.00$ $0.01$
Random $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $2.02\pm 0.04$ $-0.02\pm 0.04$ $[0.03,0.99]$
Spiral $0.01\pm 0.00$ $0.02\pm 0.00$ $1.01\pm 0.02$ $0.98\pm 0.02$ $0.01$
Then, we switch to a higher-dimensional setting, in which we test our method on correlating pairs of synthetic datasets of 4 variables. We construct a first random dataset containing 4 independently sampled variables, and then consider 3 correlation scenarios: in the first, we compute its correlation with another likewise randomly sampled dataset, hence we expect no correlation; in the second, we randomly sample 2 variables of the second dataset, while binding the other 2 to the original dataset through trigonometric functions, establishing a partial correlation between the two; finally, in the third scenario we set all variables of the second dataset to be trigonometric functions of the original variables, thus making the datasets completely correlated, although in a strongly nonlinear fashion. We compare our $I_{d}$or against Distance Correlation and Canonical Correlation Analysis (CCA) (Hotelling, 1936), which is the direct analogous of $R^{2}$ in multivariate correlation, and serves as a linear baseline. As we report in Table 2, while all methods agree in correctly finding no correlation in the first scenario, both CCA and dCor fail in the partially and completely correlated cases, while $I_{d}$or returns substantially higher correlation coefficients, showcasing its robustness to nonlinear correlations. In the Appendix (sections A.3.2 and A.3.3), we report additional results in this setting, obtained applying $I_{d}$or to differently sampled and noisy data.
[…]
Correlation $\mathbf{CCA}$ $\mathbf{dCor}$ $\mathbf{I_{d}\oplus}$ $\mathbf{I_{d}Cor}$ $\mathbf{p}$-value
(A) Absent $0.02\pm 0.01$ $00.00\pm 0$ 0.00 $8.07\pm 0.17$ $0.01\pm 0.05$ $[0.25,1.00]$
(B) Partial $0.02\pm 0.01$ $00.00\pm 0$ 0.00 $6.41\pm 0.10$ $0.35\pm 0.03$ $0.01$
(C) Complete $0.02\pm 0.00$ $0.01\pm 0.00$ $4.88\pm 0.06$ $0.67\pm 0.02$ $0.01$
4.2 A motivating example on neural representations
@@
Figure 2: Average correlation results with different methods between MNIST data and their final representations computed by a randomly initialized MLP, with variable degree of activation nonlinearity, increasing on the $x$ axis. Shaded area represents standard deviation over 10 runs with independent random initialization of MLP weights.
@@
To illustrate this phenomenon, we showcase a simple example: we consider a randomly initialized multilayer perceptron (MLP), made of $15$ fully connected layers of $784$ neurons, followed by a LeakyReLU activation. LeakyReLU is a parametric activation function, whose behavior depends on a parameter called slope: if the slope is $1$, it behaves like the identity function, rendering our MLP a linear function of the input, while lower slope values make the network nonlinear, with 00 corresponding to the standard ReLU. We feed our MLP with the MNIST (LeCun et al., 1998) dataset at increasing degrees of nonlinearity (which corresponds to decreasing the slope) and compute the correlation between the representation at the final layer and the input data, both with our method and with established baselines (SVCCA, Distance Correlation, linear kernel CKA, and RBF kernel CKA).
@@
Our method estimates a proxy of the normalized mutual information between representations, being therefore largely insensitive to nonlinear transformations of the data (as opposed to other metrics like CKA or dCor) and making it optimal to detect statistical dependencies between representations. However, this insensitivity makes it a less informative geometrical similarity index, since the two representations will have a high mutual information regardless of the nature of the nonlinear transformation.
@@ 4.3 ImageNet representations
Moving to a more realistic setting, we test our method on measuring similarity between ImageNet (Russakovsky et al., 2015) embeddings coming from different neural encoders. We consider a variety of pre-trained architectures, including CNNs (ResNet-18 (He et al., 2016) and EfficientNet-B0 (Tan & Le, 2019)), four variants of Vision Transformers (Dosovitskiy et al., 2020) (including a ViT-CNN hybrid) and three ViT-based self-supervised vision-language models (CLIP-ViT-B (Radford et al., 2021), SigLIP-ViT-B (Zhai et al., 2023) and BLIP-ViT-B (Li et al., 2022)). For all the models, we consider the output of the last hidden representation produced by the encoder, before the classification head: in supervised ViTs this choice corresponds to the last class token, while in CNNs to the pooler output. For CLIP, SigLIP and BLIP, we consider instead the image representation in the shared vision-language space.
Figure 3: Correlation results on ImageNet representations, obtained using: left Distance Correlation (dCor); right $I_{d}$or (ours). Both methods are able to detect non-negligible correlation. More baseline results are reported in the Appendix.
@@
We report the correlation results produced by our method in Fig. 3, along with the correlation scores returned by Distance Correlation (dCor), as it is the baseline method that most closely matches ours in terms of mean off-diagonal correlation ($I_{d}$or mean: $0.87$, dCor mean: $0.50$). Detailed results for SVCCA (mean: $0.46$), linear kernel CKA ($0.43$) and RBF kernel CKA ($0.44$) are reported in the Appendix in Fig. 8. In this setting, all models are computing embeddings for the same data points, hence we would expect significant correlation to be present for any given pair of models. Indeed, all methods are clearly capturing such correlation (even if with higher variance than $I_{d}$or), including SVCCA, which suggests that a relevant component of this correlation is actually linear.
@@ 4.3.1 Coarse alignment
The previous section demonstrated that our method effectively identifies strong correlations within perfectly aligned datasets of representations. We now aim to explore how the performance of the method might vary when applied to coarsely aligned data. To test this, we utilize the inherent class information of ImageNet data. Specifically, we randomly shuffle the embeddings generated by a model while keeping the labels unchanged, and then compare this modified dataset with the original dataset prior to shuffling. In other words, given a point index $i$ of class $C_{i}$, we pair it with another randomly chosen point $j$ of class $C_{j}$ with the condition that $C_{i}=C_{j}$.
This allows us to assess the robustness of our correlation estimation in less ideal conditions: we expect the correlation signal to decrease, as alignment is a crucial property for all correlation methods to detect similarities. However, as we report in Table 3, we are still able to identify correlations with high confidence, as demonstrated by the low $p$-values. The values of the correlation suggest that the number of features needed to perform the classification task is between 50 and 75%. On average, $I_{d}$or returns a correlation of $0.62$, while baselines sit in the range $0.31-0.47$. Interestingly, we observe that the models that exhibit lower correlation are CLIP, BLIP and SigLIP, all contrastive vision-language models. In two of the three cases, $I_{d}$or is even surpassed by Distance Correlation and CKA. While we have no clear explanation for such behavior, these results align with very recent results by Ciernik et al. (2024), highlighting the impact of the pre-training objective on representation similarity. For reference, we also report the $I_{d}$or results obtained when one of the two datasets is freely shuffled (irrespective of class labels). As witnessed by the high $p$-values, $I_{d}$or correctly reports a lack of correlation in such case.
Table 3: Correlation between ImageNet representations when exact alignment is broken. Results are shown, in terms of $p$-value and correlation coefficient, for the coarse alignment case, in which data are shuffled while preserving the labeling. $I_{d}$or significantly outperforms all baseline methods, with the exception of two models. The last two columns report the $I_{d}$or coefficient and $p$-value for the fully shuffled case, in which no constraint is enforced on labels, and correlation is not present, as identified by high $p$-values. All results are reported as means (excluding $p$-values, reported as ranges) over $5$ independent shufflings. Standard deviations are not reported as they are lower than 0.02 in all cases.
Model $I_{d}$or $p$-val. SVCCA CKA (lin.) CKA (RBF) dCor $I_{d}$or (rand) $p$-val. (rand)
EfficientNet $\mathbf{0.61}$ $0.01$ $0.25$ $0.17$ $0.19$ $0.25$ $-0.10$ $[0.16,0.97]$
SigLIP $0.53$ $0.01$ $0.26$ $0.58$ $\mathbf{0.59}$ $\mathbf{0.59}$ $-0.01$ $[0.09,0.86]$
ViT-B-16 $\mathbf{0.69}$ $0.01$ $0.39$ $0.34$ $0.34$ $0.44$ $0.27$ $[0.27,0.98]$
ViT-B-32 $\mathbf{0.68}$ $0.01$ $0.40$ $0.40$ $0.40$ $0.47$ $0.22$ $[0.25,0.99]$
ViT-hyb. $\mathbf{0.63}$ $0.01$ $0.39$ $0.51$ $0.52$ $0.55$ $0.27$ $[0.09,0.97]$
ViT-L $\mathbf{0.73}$ $0.01$ $0.40$ $0.45$ $0.43$ $0.55$ $0.37$ $[0.18,0.91]$
ResNet $\mathbf{0.66}$ $0.01$ $0.22$ $0.34$ $0.35$ $0.37$ $0.39$ $[0.46,0.91]$
CLIP $0.53$ $0.01$ $0.26$ $0.59$ $0.61$ $\mathbf{0.62}$ $0.21$ $[0.08,0.85]$
BLIP $\mathbf{0.50}$ $0.01$ $0.28$ $0.37$ $0.39$ $0.39$ $-0.10$ $[0.28,0.95]$
Average $\mathbf{0.62}$ $0.01$ $0.31$ $0.42$ $0.42$ $0.47$ $0.17$ $[0.08,0.99]$
@@ 4.4 Multimodal representations
We now shift our focus to a multimodal context, where we examine the similarities between hidden spaces learned by text and image encoders. We use three datasets, N24News (Wang et al., 2022), MS-COCO 2014 (Lin et al., 2014) and Flickr30k (Young et al., 2014), all consisting of image-caption pairs. This analysis will help us understand how textual and visual representations correlate when evaluated on related multimodal content. Images are encoded using a representative subset of the vision models introduced in section 4.3: two CNNs (EfficientNet-B0 and ResNet-18), two ViTs (ViT-B-16 and ViT-hybrid), and the visual branch of CLIP. For text we employ five architectures, all taken pre-trained: BERT (Devlin et al., 2019), both cased and uncased, ALBERT (Lan et al., 2019), Electra (Clark et al., 2020) and finally the text encoder of CLIP. For all text models, we consider the last representation of the class token.
We report the correlation results obtained on N24News in Fig. 4, comparing our method against Distance Correlation (dCor), which is once again the closest-performing baseline method. Our $I_{d}$or yields a mean off-diagonal correlation of $0.66$, which is noticeably higher than those of baseline methods, in the range $(0.25-0.29)$. In fact, the correlation heatmaps for all baseline methods reveal a clear block structure, as such methods are able to capture correlation only among same-modality encoders, but fail on cross-modal correlation. Full results for baseline methods are available in the Appendix (Fig. 9). Instead, our method returns significant correlation values even across different modalities, in accordance with previous findings (Maiorca et al., 2023) that proved N24News representations to be transferable across models and modalities. Experiments on Flickr30k and MS-COCO confirm the behavior observed for N24News, as discussed in the Appendix (section A.5.2). Moreover, in sections A.5.3 and A.6, we provide additional experiments on the discussed datasets to show the discriminative power of $I_{d}$or in cases with low or no correlation, while section A.7 contains additional similarity results on N24News, obtained using RTD (Barannikov et al., 2022).
Figure 4: Correlation results on N24News representations, obtained using: left Distance Correlation (dCor); right $I_{d}$or (ours). Like all other baselines we evaluate, dCor is only able to spot correlations between encoders of the same modality, while $I_{d}$or reveals significant correlation for all model pairs.
@@ 4.5 Computational resources
We performed all the computations on a NVIDIA A100 GPU, equipped with 40GB of RAM. The main computational hurdle of our method is the estimation of $I_{d}$ through TwoNN: our implementation follows closely that of Ansuini et al. (2019), which we translated to PyTorch to enable GPU acceleration. With this setup, $I_{d}$or runs in the order of $1$s on two $1024$-dimensional datasets of $30000$ points. Just like our correlation method, all the representation similarity baselines we employ greatly benefit from GPU acceleration: we used them in their PyTorch implementations provided by Miranda (2021) (SVCCA), Maiorca (2024) (CKA) and Zhen et al. (2022) (Distance Correlation), with minor adaptations. Pretrained models were obtained from the Transformers library by HuggingFace (Wolf et al., 2020), details on the checkpoints we employed are provided in the Appendix (section A.8). Our code is available at https://github.com/lorenzobasile/IDCorrelation.
@@ 5 Discussion
Our work introduces Intrinsic Dimension Correlation ($I_{d}$or), a novel and robust method for detecting complex nonlinear correlations in high-dimensional spaces. Due to its flexibility, this method can be employed in a wide range of applications, from natural language processing to computer vision and beyond (including other fields of science, like physics), offering a new type of analysis to address how machine learning models represent data. Remarkably, our results show the effectiveness of the method to detect a correlation signal in multimodal data where we know a correlation should exist but where standard methods struggle to identify any.